OpenStudy (anonymous):

1. If 9 students sharpen 18 pencils in 2 minutes, how many students will it take to sharpen 18 pencils in 1 minute? A.) 2 B.) 9 C.) 18 D.) 36 2. If y varies inversely with x, and y = –16 when x = –64, what is the constant of variation? A.) 16 B.) 64 C.) 100 D.) 1,024 3. The pair of points (3, 8) and (x, 6) are on the graph of an inverse variation. What is the missing value? A.) 4 B.) 3 C.) -1 D.) 1

OpenStudy (anonymous):

1)c

OpenStudy (anonymous):

3) d i think

OpenStudy (anonymous):

4. The weight needed to balance a lever varies inversely with the distance from the fulcrum to the weight. A 120-lb weight is placed on a lever, 5 ft from the fulcrum. What amount of weight, in pounds, should be placed 8 ft from the fulcrum to balance the lever? A.) 120 B.) 75 C.) 8 D.) 125 5. In the inverse variation equation xy/4 = -3, what is the constant of variation? A.) -3 B.) 4 C.) -12 D.) -3/4 6. What is the graph of y = 4/x?

OpenStudy (anonymous):

4) d i think

OpenStudy (anonymous):

I need help not people guessing. @skinny23

OpenStudy (anonymous):

5) c

OpenStudy (anonymous):

sorry..

OpenStudy (anonymous):

Did you just guess on all? @skinny23

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Which ones did you guess on? @skinny23

OpenStudy (anonymous):

just 3 n 4 cuz i dont know them for sure

OpenStudy (anonymous):

What about number one?

OpenStudy (anonymous):

number 1 is c

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

because if it takes 2 minutes for 9 people to sharpen 18 pencils then if you double the students for one minute

OpenStudy (anonymous):

For 2, if y varies inversely with x, then \(\Large y = {k \over x}\) where k is the constant of variation, when y = -16 x = -64 so \(\Large -16 = \frac{k}{-64}\), \(k = -16 * - 64 = 1024\) For 3, same kind of approach: y = k/x, when x = 3, y = 8 so k = 24 y = 24/x (x, 6) is on the graph y = 24/x so 6 = 24/x, x = 4

OpenStudy (anonymous):

What about 4 or 5? @Meepi

OpenStudy (help_needed_cudi):

4. The weight needed to balance a lever varies inversely with the distance from the fulcrum to the weight. A 120-lb weight is placed on a lever, 5 ft from the fulcrum. What amount of weight, in pounds, should be placed 8 ft from the fulcrum to balance the lever? (1 point) 120 75<<<<<answer 8 125

OpenStudy (help_needed_cudi):

\[120 * 5 = 8 * x \] \[8 \div 600 = 8x \div 8\]\[75 = x\]

OpenStudy (help_needed_cudi):

oh n 6. btw is a graph this is the right answer